OpenStudy (anonymous):

What is the equation of the line of symmetry for the graph of y = –2(x – 4)^2 + 3?

hero (hero):

Hi

OpenStudy (anonymous):

hi can you help me? @Hero

hero (hero):

Yes. Have you graphed this yet?

OpenStudy (anonymous):

no :/

hero (hero):

Not to worry. I graphed it for you: https://www.desmos.com/calculator/e5qkzv643d

hero (hero):

The line of symmetry occurs where two sides of the parabola on either side of the line of symmetry are mirror images of each other. The line of symmetry also intersects the vertex of the parabola.

hero (hero):

Do you know the vertex of the parabola?

OpenStudy (anonymous):

no i don't...

hero (hero):

The equation y = –2(x – 4)^2 + 3 of the graphed parabola is written in vertex form y = a(x - h)^2 + k where (h,k) is the vertex. In other words, (h,k) = (4,3) is the vertex. Which means the line of symmetry occurs where x = h.

hero (hero):

And what does h equal?

OpenStudy (anonymous):

4?

hero (hero):

Yes, correct. Thus the equation of the line of symmetry for the graph of y = -2(x - 4)^2 + 3 is x = 4

OpenStudy (anonymous):

When solving x^2 – 12x = –13 by completing the square, what value is added to both sides of the equation?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

??..

hero (hero):

You always add \(\left(\frac{b}{2}\right)^2 \) to both sides

hero (hero):

In other words, simplify \(\left(-\frac{12}{2}\right)^2\). Then add that to both sides.